Circular No. 7201

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1999 J5 (LINEAR) An apparently asteroidal object reported by LINEAR on May 12 and

17, and linked by G. V. Williams to LINEAR observations on June 8 and 10 by way of a cometlike orbit, was posted on The NEO Confirmation Page for additional observations. P. Pravec and P. Kusnirak, Ondřejov, reported that their June 12 CCD images showed a faint coma and a tail marginally visible to the southwest. Also, A. Sugie, Dynic Astronomical Observatory, reports strong condensation and a coma diameter of 12′′ on June 14. The discovery observation is given below, and additional astrometry appears on MPEC 1999-L35.

1999 UT  α_{2000}  δ_{2000}  Mag.
May 12.35863  16^{h}42^{m}19.^{s}96  −8^{o}28′01′′.6  19.0

The following orbital elements by B. G. Marsden, from 41 observations, May 12–June 14, confirm that this comet is of short period:

T = 1999 May 12.8389 TT  ω = 132.^{o}3026
e = 0.169370  Ω = 112.0017 }2000.0
q = 3.713551 AU  i = 13.7214
a = 4.470763 AU n^{o} = 0.1042633 P = 9.453 years
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 11 16^{h}23.^{m}84 − 8^{o}40.^{′}2 2.740 3.715 161.^{o}1 5.^{o}1 16.9
21 16 18.09 − 8 59.9 2.786 3.716 152.4  7.3 16.9
July 1 16 13.44 − 9 27.3 2.857 3.718 142.8  9.5 17.0
11 16 10.22 −10 01.7 2.949 3.720 133.2 11.5 17.1
21 16 08.66 −10 42.0 3.060 3.722 123.8 13.1 17.1

COMET C/1999 H1 (LEE) J. Davies and T. Kerr, Joint Astronomy Centre, report on behalf of the

collaboration noted on IAUC 7198: “Moderate resolution (R ∼ 1000–7000) 2.94–4.75-μm spectra of comet C/1999 H1 taken at the U.K. Infrared Telescope during May 20–29 UT showed strong emissions in the C-H stretch region (3.3–3.6 μm). Assuming that all of the flux in the emission feature centered near 3.51 μm (F = 4.0 × 10^{−17} W m^{−2} in a 0′′.61 × 1′′.83 aperture) is due to fluorescence from methanol, our preliminary estimate for the methanol production rate is Q(CH_{3}OH) = 2.1 × 10^{27} s^{−1}. Temporal variability was < 20 percent during the period of observations.”

1999 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green